# Convergence, Duality and Well-Posedness in Convex Bilevel Optimization

### v1

@@ Abstract
We consider the convex bilevel optimization problem, also known as simple bilevel programming. There are two challenges in solving convex bilevel optimization problems. Firstly, strong duality is not guaranteed due to the lack of Slater constraint qualification. Secondly, we demonstrate through an example that convergence of algorithms is not guaranteed even when usual subotimality gap bounds are present, due to the possibility of encountering super-optimal solutions. We show that strong duality (but not necessarily dual solvability) is exactly equivalent to ensuring correct asymptotic convergence of both inner and outer function values, and provide a simple condition that guarantees strong duality. Unfortunately, we also show that this simple condition is not sufficient to guarantee convergence to the optimal solution set. We draw connections to Levitin-Polyak well-posedness, and leverage this together with our strong duality equivalence to provide another condition that ensures convergence to the optimal solution set. We also discuss how our conditions have been implicitly present in existing algorithmic work.
keywords:
convex bilevel optimization; simple bilevel programming; duality; well-posedness
{MSCcodes}
90C25; 90C30; 90C46
@@ 1 Introduction
(1) $g^{*}=\min_{x\in X}g(x),$
where $X$ is a finite-dimensional closed convex set and $g:X\to{\mathbb{R}}$ is a convex function. Here, $g^{*}$ is the optimal value of Eq. 1. In general, problem (1) might have multiple optimal solutions, and in this case it is natural to consider a second objective to select the best solution among the optimal solutions with respect to an additional criterion. Supposing that this is given by a convex function $f:{X}\to\mathbb{R}$, this leads to the following so-called convex bilevel optimization problem (also known as simple bilevel program):
(2) $f^{*}=\min_{x\in X_{g}}f(x),\quad\text{where }X_{g}=\operatorname*{arg\,min}_{z\in X}g(z).$
@@
In contrast to a standard single-level convex optimization problem, the presence of two objective functions in the convex bilevel problem (2) brings in additional challenges, which we categorize into two groups. First, we do not have an explicit representation of the optimal set $X_{g}$ in general. This rules out the possibility of applying projection-based algorithms as well as the conditional gradient method, making both the projection onto $X_{g}$ and the linear optimization oracle over $X_{g}$ intractable. With this first challenge in mind, a stream of literature in this domain has alternatively considered the value function formulation of (2):
(3) $\min_{x\in X}\left\{f(x):\ g(x)\leq g^{*}\right\}.$
@@
1.1 Contributions
Approaches to solve single-level optimization problems cannot by default be applied to Eq. 2 (or equivalently Eq. 3), and thus it requires specific algorithms for the bilevel setting, with different analyses. In this paper, we examine the algorithmic guarantees which exist for various bilevel optimization algorithms, and provide conditions to reassure us of convergence to an optimum solution of Eq. 2. Algorithmic guarantees for Eq. 2 generally ensure that an iteratively computed sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfies the following suboptimality gap bounds
(4) $f(x_{t})\leq f^{*}+r_{f,t},\quad g(x_{t})\leq g^{*}+r_{g,t},\quad\textbf{where}\lim_{t\to\infty}r_{f,t}=\lim_{t\to\infty}r_{g,t}=0.$
Unfortunately, as we will demonstrate in Example 2.1, this is not sufficient to ensure that $\{x_{t}\}_{t\geq 1}$ converges to a solution of Eq. 2. Our study aims to analyze this carefully, and provide conditions that ensure convergence. We examine two notions of convergence: (i) convergence in objective function value Eq. 5; and (ii) convergence in distance to the optimal set Eq. 7.
Our main result establishes a previously unknown equivalence: strong duality (but not necessarily dual solvability) for the value-function formulation (3) is equivalent to ensuring convergence in objective function value. Despite the lack of Slater condition, we present two simple conditions that ensure strong duality: one based on a proximity assumption for $g$ (Section 3), and the other based on compactness of the optimal set (Section 3). While the proximity assumption is natural, it unfortunately does not ensure convergence in distance to the optimal set, as Example 3.3 demonstrates. On the other hand, we show that the compactness assumption is equivalent to two notions of Levitin-Polyak well-posedness of Eq. 2, which yields both convergence in objective value as well as in distance to the optimal set. Our strong duality equivalence is further used to provide quantifiable bounds on the distance. These findings provide algorithm-independent criteria under which iterate sequences for Eq. 2 have meaningful convergence guarantees.
1.2 Related work
@@
Our work focuses the duality theory of Eq. 2, namely the Lagrange dual of its value function formulation Eq. 3. Friedlander and Tseng [12] show that correctness of the exact regularization technique is equivalent to the existence of an optimal dual multiplier of Eq. 3, and they demonstrate a number of special cases where exact regularization holds, e.g., conic programs and exact penalty methods for convex functional constraints. Dempe et al. [8], Franke et al. [11] provide conditions to guarantee existence of an optimal dual multiplier by using more general constraint qualifications than Slater. In contrast, our work focuses on when strong duality holds regardless of whether the dual is solvable or not, and in particular we highlight that this case has important algorithmic implications for Eq. 2. Consequently, we are able to provide weaker conditions for strong duality, and thus convergence assurance, than the ones above that guarantee that an optimal dual multiplier exists. Friedlander and Tseng [12] observe that, under coercivity assumptions, strong duality holds (without solvability), which is a simple consequence of Sion’s minimax theorem, though they did not explore this further.
Finally, we note that there is a broader literature on general bilevel optimization problems with two sets of variables, which are non-convex; see Dempe [7] for details. This literature also examines optimality conditions via Lagrange multipliers, though the theory is different to the simpler convex problem Eq. 2 that we consider.
1.3 Notation and basic assumption
We let $\mathbb{R}$} and $\mathbb{N}$} denote the real and natural numbers, respectively. Given any positive integer $T$, let $T]:=\{1,2,\dots,T\}$. For any $a\in{\mathbb{R}}$, we let $a]_{+}:=\max\{0,a\}$. We define $\{a_{t}\}_{t\in[T]}$ to be the sequence of elements $(a_{1},a_{2},\dots,a_{T})$. Let $\|\cdot\|_{2}$ be the Euclidean norm. For a convex function $f:X\to{\mathbb{R}}$, we denote the gradient at $x\in X$ by $\nabla f(x)$.
We will assume that Eq. 2 satisfies the following properties throughout the paper. {assumption} The set $X$ is closed and convex, and belongs to some finite-dimensional Euclidean space. The functions $f$ and $g$ are convex and real-valued on all of $X$. Furthermore, the minimization problem Eq. 1 is solvable, i.e., $g^{*}<\infty$ and $X_{g}$ defined in Eq. 2 is non-empty and closed. Finally, the bilevel optimization problem Eq. 2 is solvable, and we denote the set of optimal solutions to Eq. 2 as
@@ 2 Super-optimality and Lagrange duality
The central issue in solving Eq. 2 is that, in practice, many algorithms very rarely find a feasible solution $\bar{x}\in X_{g}$. Instead, algorithms will generate a point $x_{t}\in X$ with $g(x_{t})>g^{*}$. Since $X_{g}\subset\left\{x\in X:g(x)\leq g(x_{t})\right\}$, the point $x_{t}$ may be super-optimal for Eq. 2, meaning that $f(x_{t})<f^{*}$. As iterative algorithms progress, the accuracy of $x_{t}$ (hopefully) improves, i.e., it is shown that $g(x_{t})\leq g^{*}+r_{g,t}$ where $r_{g,t}\geq 0$ are error terms such that $r_{g,t}\to 0$. Since $g(x_{t})\geq g^{*}$, this implies that $\lim_{t\to\infty}g(x_{t})=g^{*}$. For the outer-level objective $f$, it is often similarly shown that $f(x_{t})\leq f^{*}+r_{f,t}$, where $r_{f,t}\geq 0$ are error terms such that $r_{f,t}\to 0$. Since we might have $f(x_{t})<f^{*}$, these inequalities are only sufficient to deduce
(5) $\limsup_{t\to\infty}f(x_{t})\leq f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*},$
@@
To guard against situations like in Example 2.1, and recognizing the numerous algorithmic works which provide guarantees of the form Eq. 5, we look for conditions which ensure that whenever Eq. 5 holds, the following guarantee also holds:
(7) $\lim_{t\to\infty}f(x_{t})=f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*}.$
Another guarantee that is of interest is that the distance to the optimal solution set of Eq. 2 decreases to 0:
(8) $\lim_{t\to\infty}\operatorname{Dist}(x_{t},X^{*})=0,\quad\text{where }\operatorname{Dist}(x,X^{*}):=\inf_{z\in X^{*}}\|x-z\|_{2}.$
@@
(9) $d^{*}:=\sup_{\lambda\geq 0}q(\lambda),\quad\text{where}\quad q(\lambda):=\inf_{x\in X}\left\{f(x)+\lambda(g(x)-g^{*})\right\}.$
@@
Note that Definition 2.2 does not guarantee that the dual is solvable, but only that $q(\lambda)\to f^{*}$ as $\lambda\to\infty$. Therefore, ensuring that strong duality holds requires much weaker conditions than ensuring that the dual is solvable, as was studied in previous works [12, 8, 11].
@@
Theorem 2.3.
Consider Eq. 2 when $f,g,X$ are convex. Then the following are equivalent:
@@
    Any sequence $\{x_{t}\}_{t\geq 1}\subset X$ that satisfies Eq. 5 also satisfies Eq. 7.
@@
    Strong duality (Definition 2.2) holds for Eq. 2.
We will use the following lemma in the proof of Theorem 2.3, as well as subsequent results.
Lemma 2.4.
@@
Proof 2.5 (Proof of Lemma 2.4).
We consider two cases: $d^{*}>$-$\infty$ and $d^{*}=$-$\infty$. In the first case, there must exist some $\lambda_{0}\geq 0$ such that $q(\lambda_{0})=\inf_{x\in X}\left\{f(x)+\lambda_{0}(g(x)-g^{*})\right\}>$-$\infty$, and consequently, $q(\lambda)\geq q(\lambda_{0})>$-$\infty$ for any $\lambda\geq\lambda_{0}$. For each $t\geq 1$, choose $x_{t}\in X$ to be a point that satisfies $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})+1/t$. By definition of $q(\cdot)$, we also have $f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\geq q(\lambda_{0})$, so combining with $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})$ we get
@@
This implies that $g(x_{t})\to g^{*}$ as $t\to\infty$. Now, from the choice of $x_{t}\in X$, using the definition of $g_{*}$ we have $g(x_{t})\geq g_{*}$ and also we have
@@
The first equality follows because $g(x_{t})\to g^{*}$, and the last equality follows since $\lim_{t\to\infty}q(t+\lambda_{0})=d^{*}$ (recall the definition of $d_{*}$ and the fact that $q(\lambda)$ is non-decreasing in $\lambda$).
Now we consider the second case when $d^{*}=$-$\infty$. Then for $t\geq 1$, let $z_{t}\in X$ be a point such that $f(z_{t})+t(g(z_{t})-g^{*})\leq-t$, and define $\theta_{t}:=\frac{1}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\in[0,1]$. Let $x^{*}\in X^{*}$ be a solution to Eq. 2, and define $x_{t}:=\theta_{t}z_{t}+(1-\theta_{t})x^{*}\in X$. Notice that $g(x_{t})\leq\theta_{t}g(z_{t})+(1-\theta_{t})g(x^{*})=g^{*}+\theta_{t}(g(z_{t})-g^{*})=g^{*}+\frac{g(z_{t})-g^{*}}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\leq g^{*}+1/\sqrt{t}\to g^{*}$. Thus, $\lim_{t\to\infty}g(x_{t})=g^{*}$. On the other hand, $f(x_{t})\leq\theta_{t}f(z_{t})+(1-\theta_{t})f(x^{*})$, and by definition of $z_{t}$, we have $f(z_{t})\leq-t(1+g(z_{t})-g^{*})$, so $\theta_{t}f(z_{t})\leq-\frac{\sqrt{t}(1+g(z_{t})-g^{*})}{\max\{1,g(z_{t})-g^{*}\}}$. But this implies $\theta_{t}f(z_{t})\to$-$\infty$ as $t\to\infty$, therefore $\limsup_{t\to\infty}f(x_{t})=-\infty=d^{*}$.
Proof 2.6 (Proof of Theorem 2.3).
Suppose that strong duality holds, i.e., $f_{*}=d_{*}$. Let $\{x_{t}\}_{t\geq 1}$ be a sequence satisfying Eq. 5. Then given any $\epsilon>0$, there exists $\lambda_{\epsilon}\in[0,\infty)$ such that
@@
Suppose now that strong duality does not hold, so $d^{*}=\sup_{\lambda\geq 0}q(\lambda)<f^{*}$. Lemma 2.4 states that there exists a sequence $\{x_{t}\}_{t\geq 1}$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Since $d^{*}<f^{*}$, Eq. 5 holds for $\{x_{t}\}_{t\geq 1}$, but not Eq. 7.

### v2

@@ Abstract
We consider the convex bilevel optimization problem, also known as simple bilevel programming. Strong duality of the convex bilevel optimization problem is not guaranteed, due to the lack of Slater constraint qualification. Furthermore, convergence of algorithms is in general not guaranteed without further conditions due to possibility of super-optimal solutions. We show that strong duality (but not necessarily dual solvability) is exactly equivalent to ensuring correct asymptotic convergence of function values. We provide a simple condition that guarantees strong duality. Unfortunately, we show that this is not sufficient to guarantee convergence to the optimal solution set. We draw connections to Levitin-Polyak well-posedness, and leverage this together with our strong duality equivalence to provide another condition that ensures convergence to the optimal solution set with quantifiable bounds.
@@ 1 Introduction
$g^{*}=\min_{x\in X}g(x),$ (1)
where $X$ is a finite-dimensional closed convex set and $g:X\to{\mathbb{R}}$ is a convex function. Here, $g^{*}$ is the optimal value of Eq. 1. In general, problem (1) might have multiple optimal solutions, and in this case it is natural to consider a second objective to select the best solution among the optimal solutions with respect to an additional criterion.
Supposing that this is given by a convex function $f:{X}\to\mathbb{R}$, this leads to the following so-called convex bilevel optimization problem (also known as simple bilevel program):
$f^{*}=\min_{x\in X_{g}}f(x),\quad\text{where }X_{g}=\operatorname*{arg\,min}_{z\in X}g(z).$ (2)
@@
In contrast to a standard single-level convex optimization problem, the presence of two objective functions in the convex bilevel problem (2) brings in additional challenges, which we categorize into two groups. First, we do not have an explicit representation of the optimal set $X_{g}$ in general. This rules out the possibility of applying projection-based algorithms as well as the conditional gradient method, making both the projection onto $X_{g}$ and the linear optimization oracle over $X_{g}$ intractable.
With this first challenge in mind, a stream of literature in this domain has alternatively considered the value function formulation of (2):
$\min_{x\in X}\left\{f(x):\ g(x)\leq g^{*}\right\}.$ (3)
@@
Contributions.
Approaches to solve single-level optimization problems cannot by default be applied to Eq. 2 (or equivalently Eq. 3), and thus it requires specific algorithms for the bilevel setting, with different analyses. In this short communication, we examine the algorithmic guarantees which exist for various bilevel optimization algorithms, and provide conditions to reassure us of convergence. Algorithmic guarantees for Eq. 2 generally ensure that an iteratively computed sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfies the following inequalities:
$f(x_{t})\leq f^{*}+r_{f,t},\quad g(x_{t})\leq g^{*}+r_{g,t},\quad\textbf{where}\lim_{t\to\infty}r_{f,t}=\lim_{t\to\infty}r_{g,t}=0.$ (4)
Unfortunately, as we will demonstrate in Section 3, this is not sufficient to ensure that $\{x_{t}\}_{t\geq 1}$ converges to a solution of Eq. 2. Our study aims to analyze this carefully, and provide conditions which ensure convergence. It turns out that this can be studied by examining the regularity of the Lagrange dual of Eq. 3. As mentioned above, Eq. 3 does not satisfy Slater’s condition is by definition. That said, we can still analyze its Lagrange dual, and examine conditions for strong duality to hold. Here, we must differentiate between two cases: (i) one where the optimal dual value is equal to $f^{*}$; and a stronger case where additionally there is an optimal dual multiplier $\lambda^{*}$. We will show that the first case is in fact equivalent to ensuring that sequences satisfying Eq. 4 will in fact converge to an optimal solution of Eq. 2. This uncovers a useful hidden regularity structure in Eq. 2 that, to the best of our knowledge, has not been elaborated on previously. We conclude the paper by presenting two conditions to ensure that strong duality holds, and examine how these conditions have been implicitly present in existing algorithmic work.
Related work.
@@
Our work focuses on the duality theory of Eq. 2, namely the Lagrange dual of its value function formulation Eq. 3. Friedlander and Tseng [12] show that correctness of the exact regularization technique is equivalent to the existence of an optimal dual multiplier of Eq. 3, and they demonstrate a number of special cases where exact regularization holds, e.g., conic programs and exact penalty methods for convex functional constraints. Dempe et al. [8], Franke et al. [11] provide conditions to guarantee existence of an optimal dual multiplier by using more general constraint qualifications than Slater. In contrast, our work focuses on when strong duality holds regardless of whether the dual is solvable or not, and in particular we highlight that this case has important algorithmic implications for Eq. 2. Consequently, we are able to provide weaker conditions for strong duality, and thus convergence assurance, than the ones above that guarantee that an optimal dual multiplier exists. Friedlander and Tseng [12] observe that, under boundedness assumptions, strong duality holds (without solvability), which is a simple consequence of Sion’s minimax theorem, though they did not explore this further.
Finally, we note that there is a broader literature on general bilevel optimization problems with two sets of variables, which are non-convex; see Dempe [7] for details. This literature also examines optimality conditions via Lagrange multipliers, though the theory is different to the simpler problem Eq. 2 that we consider.
Notation and basic assumption.
We let $\mathbb{R}$} and $\mathbb{N}$} denote the real and natural numbers, respectively. Given any positive integer $T$, let $T]:=\{1,2,\dots,T\}$. For any $a\in{\mathbb{R}}$, we let $a]_{+}:=\max\{0,a\}$. We define $\{a_{t}\}_{t\in[T]}$ to be the sequence of elements $(a_{1},a_{2},\dots,a_{T})$. Let $\|\cdot\|_{2}$ be the Euclidean norm. For a convex function $f:X\to{\mathbb{R}}$, we denote the gradient at $x\in X$ by $\nabla f(x)$. We will assume that Eq. 2 satisfies the following properties throughout the paper.
Assumption 1.
The set $X$ is closed and convex, and belongs to some finite-dimensional Euclidean space. The functions $f$ and $g$ are convex and real-valued on all of $X$. Furthermore, the minimization problem Eq. 1 is solvable, i.e., $g^{*}<\infty$ and $X_{g}$ defined in Eq. 2 is non-empty and closed. Finally, the bilevel optimization problem Eq. 2 is solvable, and we denote the set of optimal solutions to Eq. 2 as
@@ 2 Super-optimality and Lagrange duality
The central issue in solving Eq. 2 is, in practice, we very rarely encounter a feasible solution $\bar{x}\in X_{g}$. Instead, algorithms will generate a point $x_{t}\in X$ with $g(x_{t})>g^{*}$. Since $X_{g}\subset\left\{x\in X:g(x)\leq g(x_{t})\right\}$, the point $x_{t}$ may be super-optimal for Eq. 2, meaning that $f(x_{t})<f^{*}$. As iterative algorithms progress, the accuracy of $x_{t}$ (hopefully) improves, i.e., it is shown that $g(x_{t})\leq g^{*}+r_{g,t}$ where $r_{g,t}\geq 0$ are error terms such that $r_{g,t}\to 0$. Since $g(x_{t})\geq g^{*}$, this implies that $\lim_{t\to\infty}g(x_{t})=g^{*}$. For the outer-level objective $f$, it is often similarly shown that $f(x_{t})\leq f^{*}+r_{f,t}$, where $r_{f,t}\geq 0$ are error terms such that $r_{f,t}\to 0$. Since we might have $f(x_{t})<f^{*}$, these inequalities are only sufficient to deduce
$\limsup_{t\to\infty}f(x_{t})\leq f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*},$ (5)
@@
To guard against situations like in Example 1, and recognizing the numerous algorithmic works which provide guarantees of the form Eq. 5, we look for conditions which ensure that whenever Eq. 5 holds, the following also holds:
$\lim_{t\to\infty}f(x_{t})=f^{*},\quad\lim_{t\to\infty}g(x_{t})=g^{*}.$ (7)
Another possible condition we consider is the that the distance to the optimal solution set of Eq. 2 decreases to 0:
$\lim_{t\to\infty}\operatorname{Dist}(x_{t},X^{*})=0,\quad\text{where }\operatorname{Dist}(x,X^{*}):=\inf_{z\in X^{*}}\|x-z\|_{2}.$ (8)
@@
$d^{*}:=\sup_{\lambda\geq 0}q(\lambda),\quad\text{where}\quad q(\lambda):=\inf_{x\in X}\left\{f(x)+\lambda(g(x)-g^{*})\right\}.$ (9)
@@
Note that Definition 1 does not guarantee that the dual is solvable, but only that $q(\lambda)\to f^{*}$ as $\lambda\to\infty$. Therefore, ensuring that strong duality holds requires much weaker conditions than ensuring that the dual is solvable, as was studied in previous works [12, 8, 11].
@@
Theorem 1.
Considering Eq. 2 when $f,g,X$ are convex, the following are equivalent:
@@
    Any sequence $\{x_{t}\}_{t\geq 1}\subset X$ that satisfies Eq. 5 also satiesfies Eq. 7.
@@
    Strong duality (Definition 1) holds for Eq. 2.
We will use the following lemma in the proof of Theorem 1, as well as subsequent results.
Lemma 1.
@@
Proof of Lemma 1.
We consider two cases: $d^{*}>$-$\infty$ and $d^{*}=\infty$. In the first case, there must exist some $\lambda_{0}\geq 0$ such that $q(\lambda_{0})=\inf_{x\in X}\left\{f(x)+\lambda_{0}(g(x)-g^{*})\right\}>$-$\infty$, and consequently, $q(\lambda)\geq q(\lambda_{0})>$-$\infty$ for any $\lambda\geq\lambda_{0}$. For each $t\geq 1$, choose $x_{t}\in X$ to be a point that satisfies $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})+1/t$. We also have $f(x_{t})+\lambda_{0}(g(x_{t})-g^{*})\geq q(\lambda_{0})$, so combining with $f(x_{t})+(t+\lambda_{0})(g(x_{t})-g^{*})\leq q(t+\lambda_{0})$ we get
@@
This implies that $g(x_{t})\to g^{*}$. Now observe that
@@
The first equality follows because $g(x_{t})\to g^{*}$, and the last equality follows since $\lim_{t\to\infty}=d^{*}$.
Now we consider the second case when $d^{*}=$-$\infty$. Then for $t\geq 1$, let $z_{t}\in X$ be a point such that $f(z_{t})+t(g(z_{t})-g^{*})\leq-t$, and define $\theta_{t}=\frac{1}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\in[0,1]$. Let $x^{*}\in X^{*}$ be a solution to Eq. 2, and define $x_{t}=\theta_{t}z_{t}+(1-\theta_{t})x^{*}\in X$. Notice that $g(x_{t})\leq\theta_{t}g(z_{t})+(1-\theta_{t})g(x^{*})=g^{*}+\theta_{t}(g(z_{t})-g^{*})=g^{*}+\frac{g(z_{t})-g^{*}}{\sqrt{t}\max\{1,g(z_{t})-g^{*}\}}\leq g^{*}+1/\sqrt{t}\to g^{*}$. Therefore $\lim_{t\to\infty}g(x_{t})=g^{*}$. On the other hand, $f(x_{t})\leq\theta_{t}f(z_{t})+(1-\theta_{t})f(x^{*})$, and by definition of $z_{t}$, we have $f(z_{t})\leq-t(1+g(z_{t})-g^{*})$, so $\theta_{t}f(z_{t})\leq-\frac{\sqrt{t}(1+g(z_{t})-g^{*})}{\max\{1,g(z_{t})-g^{*}\}}$. But this implies $\theta_{t}f(z_{t})\to$-$\infty$ as $t\to\infty$, therefore $\limsup_{t\to\infty}f(x_{t})=-\infty=d^{*}$. ∎
Proof of Theorem 1.
Suppose that strong duality holds. Let $\{x_{t}\}_{t\geq 1}$ be a sequence satisfying Eq. 5. Then given any $\epsilon>0$, there exists $\lambda_{\epsilon}\in[0,\infty)$ such that
@@
Suppose now that strong duality does not hold, so $d^{*}=\sup_{\lambda\geq 0}q(\lambda)<f^{*}$. Lemma 1 states that there exists a sequence $\{x_{t}\}_{t\geq 1}$ such that $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Since $d^{*}<f^{*}$, Eq. 5 holds for $\{x_{t}\}_{t\geq 1}$, but not Eq. 7. ∎